OpenStudy (anonymous):

find the real or imaginary solutions by factoring? x^3 + 4x^2 + 2x = -8

OpenStudy (whpalmer4):

Try factoring by grouping.\[x^3+4x^2 +2x = -8\]\[x^3+4x^2 + 2x + 8 = 0\]\[(x^3 + 4x^2) + (2x + 8) = 0\]Can you pull out any common factors from the parentheses? After doing so, do you have any ideas on how to proceed?

OpenStudy (anonymous):

2 and 1

OpenStudy (anonymous):

do i split the equation in half and have one positive and one negative solution?

OpenStudy (whpalmer4):

Show me what you get after you factor those expressions...

OpenStudy (anonymous):

(x^3 + 2x^3) = (2x +4)?

OpenStudy (whpalmer4):

No, I want you to factor \(x^3 + 4x^2\)

OpenStudy (anonymous):

x^5 + 4

OpenStudy (whpalmer4):

No, factor it! What is \[x^2(x+4)\]when multiplied out?

OpenStudy (anonymous):

x + x( x + 4)

OpenStudy (whpalmer4):

No, how do you come up with that? \[x^2(x+4) = x^2*x + x^2*4 = x^3 + 4x^2\] Does that look familiar? :-)

OpenStudy (whpalmer4):

@Relientkluver